OpenStudy (lindsey11x16):

I need help finding the indicated trigonometric value in the specified quadrant. Function: Sin theta = -3/5 Quadrant: IV Trigonometric Value: Cos theta The answer says I should have gotten 4/5 but I don't know how to get that. :/

OpenStudy (anonymous):

think of a 3 - 4 - 5 right triangle, and take the appropriate ratios

OpenStudy (anonymous):



OpenStudy (anonymous):

you can solve for \(a\) by \(a^2+3^2=5^2\) \(a^2+9=25\) \[a^2=16\]\[a=4\]

OpenStudy (lindsey11x16):

wow i was doing the math in my head WRONG the whole time i kept thinking 25-9 = 14. NO WONDER I WAS I GETTING IT WRONG. thank you!!! lol